four different point on circle whose coordinate is (mi,1/mi) then which one is ture


a. m1*m2*m3*m4=1                b.1/m1*1/m2*1/m3*1/m4 =4

Grade:

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear nishant

for simplicity let equation of circle is x2 +y2 =r2

put (m,1/m) point

so  (m)2 +(1/m)2 =r2

simplify it

  m4 -r2m2 +1=0

this equation has 4 roots m1,m2,m3,m4

product of all roots

 m1*m2*m3*m4=1


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

AskiitianExpert Pramod-IIT-R
14 years ago
Dear student Mr. Nishant for simplicity let equation of circle is x2 +y2 =r2 put (m,1/m) point so (m)2 +(1/m)2 =r2 simplify it m4 -r2m2 +1=0 this equation has 4 roots m1,m2,m3,m4 product of all roots m1*m2*m3*m4=1 Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE & AIEEE preparation. All the best. Regards, Askiitians Experts Pramod Kumar, IITR Alumni